OpenStudy (anonymous):

which solution if the equation y =4x+3?

OpenStudy (anonymous):

Please give us more information.

OpenStudy (anonymous):

(5,17) (17,5) (4,19) (4,16)

OpenStudy (anonymous):

Alright, so there's a really easy way to do this. All you have to do is plug in the numbers given in your solutions and see if the left side = right side.

OpenStudy (alexandervonhumboldt2):

subsitue (x, y) values and solve

OpenStudy (alexandervonhumboldt2):

like 17=4*5+3

OpenStudy (anonymous):

So for example, lets take (5,17), which is (x,y) 17=4(5)+3 17 = 23 so nope, this is wrong.

Nnesha (nnesha):

okay so that was your question just make sure when post question you should provide full information

Nnesha (nnesha):

hahah mod power lol good job batman

OpenStudy (anonymous):

:)

OpenStudy (jskhupmang):

whaaa you just deleted lol